40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 21 17:01:59 2022, Paco, Update, BHD, BH55 RFPD installed - part I IMG_3760.jpegIMG_3764.jpeg
    Reply  Thu Sep 22 19:51:58 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part II 
       Reply  Fri Sep 23 19:04:12 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part III 
          Reply  Thu Sep 29 18:01:14 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part IV 
             Reply  Fri Sep 30 18:30:12 2022, Anchal, Update, ASS, Model Changes 
                Reply  Wed Oct 12 14:01:23 2022, Radhika, Update, ASS, Model Changes YARM_ASS.png
                   Reply  Wed Oct 12 18:18:16 2022, rana, Update, ASS, Model Changes 
    Reply  Fri Sep 23 14:10:19 2022, Radhika, Update, BHD, BH55 RFPD installed - part I IMG_3766.jpegIMG_3770.jpegIMG_3773.jpeg
       Reply  Fri Sep 23 18:31:46 2022, rana, Update, BHD, BH55 RFPD installed - part I 
          Reply  Mon Sep 26 11:39:37 2022, Paco, Update, BHD, BH55 RFPD installed - part II PXL_20220926_175010061.MOTION.jpgBH55_Transimpedance_Measurement.pdf
             Reply  Thu Oct 6 11:12:14 2022, Anchal, Update, BHD, BH55 RFPD installation complete 
Message ID: 17150     Entry time: Wed Sep 21 17:01:59 2022     Reply to this: 17152   17155
Author: Paco 
Type: Update 
Category: BHD 
Subject: BH55 RFPD installed - part I 

[Radhika, Paco]

Optical path setup

We realized the DCPD - B beam path was already using a 95:5 beamsplitter to steer the beam, so we are repurposing the 5% pickoff for a 55 MHz RFPD. For the RFPD we are using a gold RFPD labeled "POP55 (POY55)" which was on the large optical table near the vertex. We have decided to test this in-situ because the PD test setup is currently offline.

Radhika used a Y1-1025-45S mirror to steer the B-beam path into the RFPD, but a lens should be added next in the path to focus the beam spot into the PD sensitive area. The current path is illustrated by Attachment #1.

We removed some unused OPLEV optics to make room for the RFPD box, and these were moved to the optics cabinet along Y-arm [Attachment #2].

 


[Anchal, Yehonathan]

PD interfacing and connections

In parallel to setting up the optical path configuration in the ITMY table, we repurposed a DB15 cable from a PD interface board in the LSC rack to the RFPD in question. Then, an SMA cable was routed from the RFPD RF output to an "UNUSED" I&Q demod board on the LSC rack. Lucky us, we also found a terminated REFL55 LO port, so we can draw our demod LO from there. There are a couple (14,15,20,21) ADC free inputs after the WF2 and WF3 whitening filter interfaces.


Next steps

  • Finish alignment of BH55 beam to RFPD
  • Test RF output of RFPD once powered
  • Modify LSC model, rebuild and restart
Attachment 1: IMG_3760.jpeg  547 kB  Uploaded Wed Sep 21 18:26:03 2022  | Hide | Hide all
IMG_3760.jpeg
Attachment 2: IMG_3764.jpeg  216 kB  Uploaded Wed Sep 21 18:26:13 2022  | Hide | Hide all
IMG_3764.jpeg
ELOG V3.1.3-